OpenStudy (anonymous):

what is 2 and 2/3 divided by 5/6 and how to you find the answer

OpenStudy (mertsj):

Can you change 2 2/3 to an improper fraction?

OpenStudy (jdoe0001):

recall that \(\bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (anonymous):

2 and 2/3 = 8/3. When you divide fractions you take take the reciprocal of the dividend and multiply it with the number you are dividing with. For example with your problem 2 and 2/3 divided by 5/6, you first change 2 and 2/3 to an improper fraction which is 8/3 then multiply it with the reciprocal of the dividend which is 6/5. The reciprocal is basically the fraction flipped. So now just do 8/3 times 6/5. Now find the answer!